OpenStudy (anonymous):

what is the value of n in this question 3(8n-1)+8(n+4)=40

OpenStudy (anonymous):

3(8n-1)+8(n+4)=40 (24n-3)+(8n+32)=40 32n+29=40 32n=11 n=11/32 or 0.34375

OpenStudy (anonymous):

thanks